A 56-centimeter-long board is cut into 3 pieces. The longest piece is twice as long as the shortest piece. The other piece piece is 4 cm longer that the shortest piece. How long is each piece?
Answer by Cromlix(4381)

Shortest piece = x
Longest piece = 2x
Other piece = x + 4
x + 2x + x + 4 = 56
4x + 4 = 56
4x = 56 - 4
4x = 52
x = 13cm
Shortest piece = 13cm
Other piece = 17cm
Longest piece = 26cm
Total = 56cm.
Hope this helps.
:-)